Electronic Journal: Southwest Journal of Pure and Applied Mathematics Internet: http://rattler.cameron.edu/swjpam.html ISSN 1083-0464 Vol. No. 01, 1996, pp. 51–61 Submitted: Feb 11, 1996. Published: June 15, 1996.

COMMON FIXED POINTS FOR δ-COMPATIBLE MAPPINGS

R.A.RASHWAN AND M.A.AHMED

Abstract. In this paper, using the condition of δ-compatibility between a set-valued mapping and a singlevalued mappin g due to Jungck and Rhoades [6], we generalize a common fixed point theorem of K han and Kubiaczyk [7] in a complete metric space. Their Theorem extends well kno wn results due to Chang [1], Khan et al. [8], and Singh-Whitfield [13]. We give an example to support our extension.

(1991) A.M.S. Subject Classification Codes. R54H25.

Key Words and Phrases. Common fixed points, δ-compatibility mappings, complete metric spaces.

Department of mathematics, Faculty of science, Assiut University, Assiut. EGYPT E-mail Address: assiut@frcu.enu.eg \begin{matrix}c \\ ◯ \end{matrix}1996 Cameron University

Typeset by AMS-TEX